# Pulling Back the Curtain on ReLU Networks

### v3

@@ 5.2 The implicit kernel machine and the faithfulness of excitation pullbacks
Recall that $\omega^{(t)}=\Omega((y^{(t)},\theta^{(t)}))\in\mathbb{R}^{P_{0}}$ (Equation 16) encodes the network’s parameters at time $t$ and that:
$f^{(t)}(x)$ $=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle$
$f^{(t)}_{\pi}(x)$ $=\langle\omega^{(t)},\phi_{\pi}(x)\rangle$
If the Hypothesis 1 holds, then, starting from time $t_{\pi}$ the training dynamics of $\omega^{(t)}$ closely resembles the training dynamics of the kernel machine given by $f^{(t)}_{\pi}$. In fact, $(\phi_{\tilde{G}^{(t)}}-\phi_{\pi})$ can be interpreted as a feature noise added to the fixed feature map $\phi_{\pi}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$ at time $t$ of training.
Note 5.1 Hypothesis 1 implies that gradient descent actually learns the implicit kernel machine $f_{\pi}$ and the model $M$ is just a noisy wrapper that allows us to feasibly approximate the extremely-high-dimensional feature function $\phi_{\pi}$.
Since $\pi$ is assumed to be piecewise constant, then also it’s pullback $v_{\pi}$ is piecewise constant and, by Equation 15, we have
@@
This means that the pullback $v_{\pi}^{\top}$ can actually be interpreted as the gradient of the kernel machine $f_{\pi}$. We can feasibly approximate the direction of that gradient by $v_{\Gamma}$, which is a better approximation than that of $v_{G}$ and therefore $v_{\Gamma}^{\top}$ is actually more faithful to the network’s decision than the standard gradient (if the Hypothesis 1 holds).
Also notice that in order to compute the path space dot product $\langle\phi_{\Lambda}(x),\phi_{\Lambda}(x^{\prime})\rangle$ it suffices to compute the following product kernel:
@@
In particular, we may feasibly study the properties of $\langle\phi_{\pi}(x),\phi_{\pi}(x^{\prime})\rangle$ with the help of $K_{\Gamma}$ kernel. Notice that $K_{\Gamma}$ encodes the idea behind Deep Features and thus offers a possible explanation for why the latter is an effective perceptual metric (Zhang et al., 2018), as it basically expresses the path-space dot product under the feature map $\phi_{\pi}$ (up to the scaling), tapping into the internal representation of $f_{\pi}$.
@@ 5.3 The effectiveness of Batch Normalization
Batch Normalization (Ioffe and Szegedy, 2015) has become one of the most influential architectural innovations in modern deep learning, largely due to its ability to dramatically accelerate training and improve generalization. While the standard explanation attributes these benefits to reduced internal covariate shift and smoother optimization landscapes (Santurkar et al., 2019), our framework offers an alternative and complementary perspective.
Note that BN basically removes the bias from inputs and largely prevents the emergence of data-dominant output neurons, enforcing their uniform activation across the batch. This facilitates both the neuron specialization and more uniform usage of the layer width. Highly activated neurons are therefore considerably better aligned (directionally) with the particular input then the less activated ones and have larger impact on the weight update during training, which reinforces their role. All these phenomena favor quicker stabilization of path activations, i.e. faster emergence of the path tensor feature and the corresponding linear structure of network representations, improving the training speed and representational quality.
Interestingly, this resonates with the recent line of work on B-cos networks (Böhle et al., 2022), which explicitly promote weight-input alignment as a mechanism for interpretability. The alignment of excitation pullbacks suggests that the role of B-cos transform is already performed by the BN layer in standard architectures, but this is obfuscated by the pullbacks of weakly excited neurons contaminating the gradient, see Section 4.1.
5.4 ReLU network as content-addressable storage
@@
Under the path stability hypothesis, ReLU networks can be interpreted as an associative memory system with excitation pullbacks as its content. Recall the Note 3.1 where we interpret the pullback vector field $v_{G}$ as binary filtering of atoms. This can be understood as computing the target location $\tilde{G}(x)\in\mathbb{R}^{\mathcal{P}_{1}}$ in the network’s memory space $\mathbb{R}^{\mathcal{P}_{1}}$. During training, for a positive example $x$, the network updates its weights to pull $v_{G}^{(t)}(x)$ closer to $x$ (due to Equation 7), i.e. it combines $x$ with the value $v_{G^{(t)}}(x)$ stored at the location $\tilde{G}^{(t)}(x)$. Under the path stability condition this location remains largely the same, oscillating around $\pi(x)$ and therefore $v_{\pi}(x)$ (and $v_{\Gamma}(x)$) contains the most coherent information. This is validated by the perceptual alignment of excitation pullbacks.
@@
During inference, the network first computes the location $\tilde{G}(x)$ and then takes the dot product of $x$ with the stored memories $v_{(c,G)}(x)$ for every class $c$, returning the class of the best-matching memory. According to the path stability hypothesis, we can extract the actual memories by computing $v_{(c,\Gamma)}(x)$ instead (the excitation pullback for class $c$). It may however change the network’s prediction if the pullbacks weren’t used during the training.
@@ 6 Empirical validation
This section describes how we visualize the excitation pullbacks. To this end, we focus on popular and representative classes of ReLU-based architectures pretrained on ImageNet and available via the torchvision library, such that it is straightforward to replace most or all occurrences of ReLU and MaxPool2d (see Section 6.3) layers with our alternative variants, using a single recursive function over the child modules. Based on these criteria, we selected ResNet50, VGG11_BN, and DenseNet121.
Notably, determining the optimal choice of $\sigma_{\ell}$ for every neuron or layer in a given network is outside the scope of this work. Instead, we seek a single global set of hyperparameters that performs reasonably well across all layers of a given model - and ideally, across all selected architectures. Our goal is to evaluate whether the resulting excitation pullbacks indeed exhibit a substantially improved perceptual alignment compared to standard gradients.
@@ 6.2 Technical details
We use resnet50, vgg11_bn and densenet121 models with the flag pretrained=True from the torchvision.models library.555see Section 7.3 for more context. Gradients are visualised by torchvision.utils.make_grid(scale_each=True). We seed the dataloader with torch.Generator().manual_seed(314).
@@ 7.1 Related work
In (Linse et al., 2024) it is shown that setting the negative slope of LeakyReLU considerably higher in the backward pass than the forward pass leads to much better aligned global feature vizualisations via activation maximization, which can be directly explained by our results as LeakyReLU is equivalent to setting constant excitation function for each of the real half-lines (e.g. 1 for positive values and 0.3 for non-positive). In (Horuz et al., 2025) they generalize this idea to ReLU’s with different surrogate gradients, which is even closer to our approach, indicating that particular selection of surrogate ReLU gradient (e.g. B-SiLU) consistently improves generalization performance. Combined with our findings this leads to the open question, whether computing excitation pullbacks (instead of vanilla gradients) during training can improve generalization, and to what extent. Intuitively, one should not differentiate through the excitation gates explicitly as they may saturate and overfit; nevertheless, they can still be computed in the backward pass as alternative gradients.
@@
7.2 On the possible biological plausibility of the implicit linear model
Intuitively, the relative stability of the gating mechanism throughout training is reminiscent of the approximately fixed synaptic connections observed in biological neural systems. In this analogy, the discrete gating pattern can be thought of as defining a persistent connectivity structure - akin to neural connections - while the excitation function describes the level of activity (excitation) of each neuron, resembling the frequency-based coding of information.
One may even venture a more speculative interpretation: the excitation pullbacks could be loosely associated with internal subjective representations, i.e. qualia - subjective, localized manifestations of experience that emerge from the brain activity taken as a whole - in a way similar to how excitation pullbacks emerge from the gating patterns in ReLU networks. It makes even more sense in the light of Section 5.4, where we interpret excitation pullbacks as the content of the network’s associative memory. This may open exciting direction in neuroscience and the philosophy of mind.
@@ 7.3 Limitations
Our experiments were run on pretrained torchvision models downloaded with the flag pretrained=True. This flag is now deprecated in the recent versions of the library, in particular it returns the older ResNet50_Weights.IMAGENET1K_V1. We’ve observed that it’s harder (but still possible) to achieve similar quality of excitation pullbacks for the more recent ResNet50_Weights.IMAGENET1K_V2 weights, at least when using a single global $\sigma$ for all the neurons, which, as we’ve indicated earlier, might by an oversimplified approach. We speculate that this is because of different distribution of pre-activations in newer versions of the pretrained models, which have been heavily regularized.
The formal theory developed in this paper concerns specifically the ReLU networks. Intuitively, similar arguments should hold for other activation functions, in particular the ones similar to ReLU, e.g. SiLU and GELU. Initial experiments for those architectures show promise but require certain refinement of the computation of $\Gamma$ to account for the more complex gradients. We haven’t yet explored the potential applicability of our method to attention-based models.

### v4

@@ 5.2 The implicit kernel machine and the faithfulness of excitation pullbacks
Whenever the Hypothesis 1 holds, then, starting from time $t_{\pi}$, the network becomes highly positively correlated with the kernel machine $f^{(t)}_{\pi}$ that operates on the same feature space and acts via the same vector of weights $\omega^{(t)}\in\mathbb{R}^{\mathcal{P}_{0}}$ as the network. In fact, $(\phi_{\tilde{G}^{(t)}}-\phi_{\pi})(x)=(\tilde{G}^{(t)}(x)-\pi(x))\otimes x$ can be interpreted as a feature noise added to the fixed feature map $\phi_{\pi}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$ at time $t$ of training. Thus, $M$ can itself be considered a noisy kernel machine. This offers an interesting angle to study the training dynamics of neural networks. Additionally, since $\pi$ is assumed to be piecewise constant, then so is it’s pullback vector field $v_{\pi}$ and, by Equation 15:
@@
This means that $v_{\pi}^{\top}$ can be interpreted as the gradient of the kernel machine $f_{\pi}$. Since we have $\tilde{\Gamma}(x)\sim_{prop}\pi(x)$ by Hypothesis 1, then also $v_{\Gamma}\sim_{prop}v_{\pi}$ by Equation 14, meaning that we can directionally approximate that gradient with excitation pullback, i.e. the kernel machine $f_{\pi}$ is directly computable. Additionally, the covector $v_{\Gamma}^{\top}$ can be considered as even more faithful to the network’s decision than the standard gradient, establishing a reliable explanation method (depending on the degree to which the path stability condition is satisfied).
Notice that, in order to find the path space dot product $\langle\phi_{\Lambda}(x),\phi_{\Lambda}(x^{\prime})\rangle$, we can compute the following product kernel:
@@
In particular, $K_{\Gamma}(x,x^{\prime})=\langle\phi_{\Gamma}(x),\phi_{\Gamma}(x^{\prime})\rangle=\langle\tilde{\Gamma}(x)\otimes x,\tilde{\Gamma}(x^{\prime})\otimes x^{\prime}\rangle\approx\langle\mu\pi(x)\otimes x,\mu^{\prime}\pi(x^{\prime})\otimes x^{\prime}\rangle=\mu\mu^{\prime}\langle\phi_{\pi}(x),\phi_{\pi}(x^{\prime})\rangle$. Assuming $\mu,\mu^{\prime}\approx 1$ this lets us tap into the path space dot product under the feature map $\phi_{\pi}$, positioning $K_{\Gamma}$ as a strong candidate for the measure of similarity. Incidentally, $K_{\Gamma}$ is equivalent to Deep Features, which is already empirically confirmed to be an effective perceptual metric (Zhang et al., 2018).
@@ 5.3 The effectiveness of Batch Normalization
Batch Normalization (BN) (Ioffe and Szegedy, 2015) has become one of the most influential architectural innovations in modern deep learning, largely due to its ability to dramatically accelerate training and improve generalization. While the standard explanation attributes these benefits to reduced internal covariate shift and smoother optimization landscapes (Santurkar et al., 2019), our framework offers an alternative and complementary perspective. We even suspect that BN might be crucial for the early emergence of the path tensor feature. We sketch the arguments below, assuming that the affine parameters of BN are close to their default values (i.e. at the initialization or with the affine=False flag).
The normalization performed by BN nullifies the input bias and enforces neuron specialization, as any specific unit after BN gets highly excited only for the fraction of inputs it is most aligned with (by dot product). Additionally, it forces every individual neuron to be inactive for around half of the inputs. Assuming low intra-layer correlation of weights, this encourages the network to activate a wide range of paths across the batch and only a relatively small set of paths per input, meaning that the tensors $\{\tilde{G}^{(t)}(x):x\in X\}\subset\{0,1\}^{\mathcal{P}_{1}}$ are sparse and, together, they cover a wide range of coordinates. This property is preserved by the tensor product and therefore transferred to the image $\phi_{\tilde{G}^{(t)}}(X)=\{\phi_{\tilde{G}^{(t)}}(x):x\in X\}\subset\mathbb{R}^{\mathcal{P}_{0}}$ under the network feature function $\phi_{\tilde{G}^{(t)}}$ (see Equation 23). This facilitates the linear separability of $\phi_{\tilde{G}^{(t)}}(X)$, and, in general, should significantly benefit convergence. Since we have not invoked Hypothesis 1 here, this reasoning can be viewed as supporting evidence in its favor.
5.4 Classification ReLU network as content-addressable storage
@@
Hypothesis 1 allows us to view ReLU classification network as an associative memory with excitation pullbacks as its content. Recall the Note 3.1 where we interpret the pullback vector field $v_{G}$ as a binary filtering of atoms. This can be understood as computing the target location $\tilde{G}(x)\in\mathbb{R}^{\mathcal{P}_{1}}$ in the network’s “memory space” $\mathbb{R}^{\mathcal{P}_{1}}$. Given a positive (belonging to our fixed class $c$) example $x$ during training, the network updates its weights to pull $v_{G}^{(t)}(x)$ closer to $x$ (in dot product, see Equation 7), i.e. it combines $x$ with the value $v_{G^{(t)}}(x)$ stored at the memory location $\tilde{G}^{(t)}(x)$. Under the path stability condition this location remains largely the same, oscillating around $\pi(x)$. Moreover, Batch Normalization encourages the locations to be well-separated, as argued in Section 5.3. Therefore, we can expect $v_{\pi}(x)$ to store coherent features extracted from the training data distribution, which seems to be validated by the perceptual alignment of excitation pullbacks $v_{\Gamma}$.
@@
During inference, the network first computes the memory location $\tilde{G}(x)$ and then takes the dot product of $x$ with the stored features $v_{(c,G)}(x)$ for every class $c$, returning the class of the best-matching memory. According to the path stability hypothesis, we can extract the actual, coherent memories by computing $v_{(c,\Gamma)}(x)$ instead (the excitation pullback for class $c$). This may however change the prediction, suggesting that excitation pullbacks should be used during training in place of vanilla gradients, see Section 7.1.
@@ 6 Empirical validation
This section describes how we visualize the excitation pullbacks. We focus on popular and representative classes of ReLU-based architectures pretrained on ImageNet and available via the torchvision library, such that it is straightforward to replace most or all occurrences of ReLU and MaxPool2d (see Section 6.3) layers with our alternative variants, using a single recursive function over the child modules. Following these criteria, we selected ResNet50, VGG11_BN, and DenseNet121, though we obtained encouraging preliminary results on a broader range of architectures, including SiLU and GELU based.
Note that determining the optimal choice of $\sigma_{\ell,i}$ for each neuron in a given network is left outside the scope of this work. This is because a single global $\sigma$ performed reasonably well across all three selected architectures, indicating significant robustness of the method to the hyperparameter choice. Nevertheless, we expect that tuning neuron-specific $\sigma_{\ell,i}$ functions could substantially improve the quality of feature attributions.
@@ 6.2 Technical details
We use resnet50, vgg11_bn and densenet121 models with the flag pretrained=True from the torchvision.models library (see Section 7.2). Gradients are visualised by torchvision.utils.make_grid(scale_each=True). We seed the dataloader with torch.Generator().manual_seed(314).
@@ 7.1 Related work
In (Linse et al., 2024) it is shown that setting the negative slope of LeakyReLU considerably higher in the backward pass than the forward pass leads to much better aligned global feature vizualisations via activation maximization, which can be directly explained by our results as LeakyReLU is equivalent to setting constant excitation function for each of the real half-lines (e.g. 1 for positive values and 0.3 for non-positive). In (Horuz et al., 2025) they generalize this idea to ReLU’s with different surrogate gradients, which is even closer to our approach, indicating that particular selection of surrogate ReLU gradient (e.g. B-SiLU) consistently improves generalization performance. Combined with our findings this leads to the question, whether computing excitation pullbacks (instead of vanilla gradients) during training can improve both generalization and interpretability, including other activations. Intuitively, one should not differentiate through the excitation gates explicitly as they may saturate and overfit; nevertheless, they can still be computed in the backward pass as alternative gradients.
@@
Our method resonates with the recent line of work on B-cos networks (Böhle et al., 2022), which explicitly promotes weight-input alignment as a mechanism for interpretability. The perceptual alignment of excitation pullbacks suggests that the role of B-cos transform is already performed by the BN layer, which aligns highly excited neurons with inputs, but this is obfuscated by the pullbacks of weakly excited neurons contaminating the gradients (see Section 4.1).
@@ 7.2 Limitations
Our experiments were run on pretrained torchvision models downloaded with the flag pretrained=True. This flag is now deprecated in the recent versions of the library, in particular it returns the older ResNet50_Weights.IMAGENET1K_V1. We’ve observed that it’s harder (but still possible) to achieve similar quality of excitation pullbacks for the more recent ResNet50_Weights.IMAGENET1K_V2 weights. We speculate that this is because of different distribution of pre-activations in newer, heavily regularized versions of pretrained models.
The formal theory developed in this paper concerns specifically the ReLU networks. Intuitively, similar arguments should hold for other activation functions, in particular the ones similar to ReLU, e.g. SiLU and GELU. Initial experiments for those architectures show promise but require modifying $\Gamma$ to account for the more complex gradients. We haven’t yet explored the potential applicability of our method to attention-based models.
@@
A key limitation of our work is that the contributions remain preliminary and partly speculative. A more rigorous formal treatment and supporting empirical studies are required, which we defer to future work.